T-Mobile 11:14 AM X Problem Set 2_Spring 2018.pdf 1- One gram of liquid bemzene (CH)is bumed in a bomb calorimeter. The temperature before ignition was 20826 C,and the temperature afher the combustion was 25.000 C. This was an adiabatic calorimeter (constant volame). The heat capacity of the bomb, the water around it, and the contents of the bomb before the combustion was 10,000 JK. Calculate the ??-foe C.H.1) at 298 15 K teom these data. Assume that the water produced in the combustion is in the liquid stale and the carbon dioxide produced in de combustion is in the gas state. You waill need the balanced reaction for the combustion of bennene (5 pts) 2- Superoooled water is liquid waler that has been cooled below its normal freering point. This stae is thermodymamically unstable and converts to ice isreversibly. If I mol of supercooled waler are transformed into ice at T-263.0 K and at a constant peessure of 1.00 atm, calculate the enthalpy change ( and entropy change (AS) of the waler, ASg, and AS3 pts) 3- Calculale the entropy change when a. Two mole of H:Og) are cooled irreversibly at constant pressure from 120C to 100C.(2 pts) b. One mole of H:Og) is expanded at constant pressure of 2 bar from an oripinal volume of 20 L to a final volume of 25 L You can consider the gas to be idcal. (3 pts) c. One hundred grams of H:O(s) at-1C and 1 bar are heated to H:Ol) at 10C and I bar. (3 pts) 4 Consider the reaction that converts pyrnic acid (C COCOO) into gasoous acctaldchyde (CCHO) and gaseous COs. which is catalysed in aqueous solution by the enzyme pyruvate decarboxylase. Fee pynuvic acid use the themodynamic values for the liquid state. 5. Consider a systern that has four separale compartments containing 1.0 mol of He, 3.0 mol ofNe, 2.0 mol of Ar, and 2.5 mol of Xe at 298.15 K. The pressure in each compartment is 1 bar. Calllate AS(2 pes 6. Calculate the Aena for the reactiot, 6COig) + 6H:00, ? CA{i:0 (s) + 60a(g) at T-310K a. Use AG-AM-TAS3 ps) =m@-n)(3 pts, b. Use -Froen a Differential Scanning Calorimetry (DSC) esperiment int is determined that the denaturation of a protein results in a change in enthalpy(Mis 640.1 kJ mot at a melting tmperature T-340 K at a pressure of 1 atm a. Calculate the AS for this denaturation at T-340 K and P- 1atm. (2 pts) b. Determine the entropy and enthalpy changes at 7- 310 K assuming AG-837JK moli independent of the temperature between T-330 and T-340Kpts) c. Calculate the Gibbs energy change (AG) for the protein denaturation at T- 310 K and at T-340 K(3 pts) 8- Calculate the maximum work that can be done by a reversible engine operating between 500 and 200 K if 1000 J is absorbed at S00 K 2 pts) Open With Print Please answer all, this is for physical chemistry. Thank you
Answer #1

1)

Given data,

T1=208260C +273 =21099K   

T2=250000C +273 = 252730 K

\DeltaT = T2-T1

=252730-21099

= 4174K

CALCULATION OF \DeltaE ;

\DeltaE= -CV\DeltaT Cv=10000 j/k is given     \DeltaT=4174 K

  \DeltaE= -CV\DeltaT

= -10000x4174

= 4174x104 J

Calculation of standard entholpy \DeltaH=\DeltaE + \Deltan(RT)

                                               \DeltaE= 41740 Kj         \Deltan=-3/2 (6-15/2)    T=298.15K

                                              \DeltaH= -41740 + (-3/2)0.0821X298.15

                                               =- 41740   +   (-0.123)X298.15

                                              =-41740-36.672

                                        \DeltafH 0 = - 41776.672 KJ